Help! How do I solve this??
(x-1)(x+3)=5

Answer by fetter6(6)

Solve (x-1)(x+3)=5.
The best way to do this is to set the equation to 0. So subtract 5 from both sides to give
(x-1)(x+3)-5=0
Now multiple out (x-1)(x+3) to give
(x^2+3x-x-3)-5=0 => x^2+2x-8=0.
Now factor the left hand side:
x^2+2x-8=(x+4)(x-2)=0 => x=-4,2
Your solution is x=-4 or x=2 QED